OpenStudy (anonymous):

What is the simplified form of 5^3 x 5^3?

OpenStudy (thomas5267):

\[ \begin{align*} &5^3\times5^3 \\ =&5^{3+3} \\ =&5^6 \\ =&15625 \end{align*} \]

OpenStudy (anonymous):

thank you :)

OpenStudy (thomas5267):

Give me a medal then!

OpenStudy (anonymous):

Dat medal greed :D

OpenStudy (thomas5267):

I am just hoping for a 60+ SmartScore...